OpenStudy (deadshot):

What is the focus of the parabola given by the equation x = 1 over 44 y^2?

OpenStudy (nathan917):

The vertex is (0,0) and the parabola opens right. 1/44= 1/(4p) P=11 The focus is 11 units right of the vertex, (11,0)

OpenStudy (nathan917):

@DeadShot

OpenStudy (nathan917):

Hey @amistre64 Has open study Been doing Updates a lot?

OpenStudy (amistre64):

lol, i wish. theyve only been crashing alot as far as i can tell.

OpenStudy (nathan917):

It goes off every Day.